# Structure-aware World Model for Probe Guidance via Large-scale Self-supervised Pre-train

### v1

@@ Abstract
With the rise of deep learning [4, 6, 7, 21], AI technology has shown great potential in improving the efficiency of echocardiography analysis. For example, Ouyang et al. [14] proposed a video-based deep learning algorithm, i.e., EchoNet-Dynamic, that can automatically make accurate assessments of cardiac function. More importantly, with the rise of large language models [2, 16, 17, 20] and multi-modal learning [8, 10, 11, 15], the interpretation of echocardiograms [3] has shown increasing improvement, demonstrating excellent performance in tasks such as pulmonary artery pressure estimation, left ventricular hypertrophy, heart failure, and left atrial enlargement. These AI-assisted diagnostic tools have demonstrated performance almost comparable to human experts. However, this comes with a prerequisite: the acquisition of high-quality echocardiograms. In regions with scarce medical resources, there are often no sonographers available to obtain high-quality echocardiograms. In such cases, the powerful capabilities of these AI-assisted diagnostic tools cannot be fully utilized. Few works [5, 9, 13] have focused on how to use AI technology to assist inexperienced sonographers in accurately acquiring target planes. Recently, Jiang et al. [9] proposed a cardiac dreamer, which only focuses on supervised learning of the 3D-structure of the heart and serves as a "heart map" for the probe guidance task. This work demonstrates a potential AI-assisted scanning method that is expected to improve the scanning skills of novices.
@@
According to the clinical experience of doctors, understanding both the 2D and 3D structures is crucial for efficient scanning. For example, when you need to adjust the probe’s viewing angle on a particular plane to capture specific anatomical structures, you must have a good understanding of the spatial positions of those anatomical structures on that plane (Fig.1 Left). Thus, in this paper, we propose a 2D-3D joint structure-aware pre-training framework to obtain a data-driven cardiac world model that benefits ultrasound scanning. Specifically, we require the world model to learn important spatial relationships in the following ways: (1) for understanding two-dimensional structures, we use a masking approach that requires the world model to predict features at adjacent positions in the two-dimensional plane; (2) for understanding three-dimensional structures, we provide information on the positional changes of two planes in 3D space, requiring the world model to predict the features of the target plane after the positional change. We further collected expert operational data on acquiring the most common ten standard planes from 364 routine clinical scans, resulting in 1.36 million sample pairs gathered by three certified sonographers, to enhance the model’s learning of generalizable cardiac structure knowledge. Results on the downstream probe guidance tasks indicate that the proposed structure-aware pre-training method learns useful knowledge for assisting in the acquisition of echocardiograms.
@@ 2.1 Preliminary
Context and Prediction. For the context features, patches not belonging to the target blocks are randomly selected and input into the context encoder to obtain features . Subsequently, the predictor (world model) generates the features of target blocks based on the context features and positional embedding which indicate the location of context and target blocks. We denote the predicted target blocks’ feature as .
@@
Echocardiography involves acquiring high-quality echocardiogram and subsequently conducting analysis and diagnosis based on these images. However, previous researches [3, 14] have primarily focused on understanding and analyzing the two-dimensional plane, neglecting how AI can assit in acquiring high-quality echocardiogram. While understanding the 2D structure of a plane can aid in scanning, especially when some features of the standard view are already visible, this alone is insufficient. Particularly when transitioning from one standard view to another, a deep understanding of the heart’s 3D structure is essential. Therefore, we propose a 2D-3D joint structure-aware pre-training framework, as illustrated in Fig.1 (a). The core insight is predicting the visual features of structures at target locations based on given 2D and 3D positional conditions, thereby learning the mapping relationship between spatial positions and visual features. Next, we provide details of the modeling and training method.
@@ 3.1 Implementation Details
Downstream task. For the probe guidance task, we adopted the framework and procedure proposed in [9], as shown in Fig.2 (b). The input for this task is an ultrasound image, and the output is the probe position adjustment needed to achieve a specific standard view. This task aims to assist junior ultrasound medical personnel in scanning, enhancing the success rate and quality of view acquisition. During fine-tuning, our pre-trained world model was loaded and optimized for 5 epochs with a batch size of 1024 on 8 Nvidia RTX-4090 GPUs. The learning rate was set to 1e-4, using a cosine scheduler with a final learning rate of 1e-6. The optimizer was set to AdamW.
@@ 4 Conclusion and Discussion
In this work, we propose a 2D-3D joint structure-aware pre-training framework to enhance the cardiac world model’s understanding of spatial relationships within two-dimensional structures on a single view and the three-dimensional spatial relationships between different views. We innovatively designed a self-supervised learning task that predicts visual features based on both two-dimensional and three-dimensional spatial information. To support large-scale self-supervised pre-training, we collected over a million ultrasound image and 3D pose data pairs. After pre-training on the large-scale dataset, considerable improvement were observed in downstream probe guidance tasks across the ten standard views.. In the future, we will attempt to validate our probe guidance model in real clinical settings, aiming to directly translate algorithmic improvements into enhanced medical outcomes or increased efficiency. Additionally, this algorithm has the potential to serve as the decision-making brain for autonomous ultrasound scanning robots, promoting the realization of fully autonomous cardiac scanning.
@@
This work was supported in part by the National Key R&D Program of China (2021ZD0140407) and the NSFC (62321005).

### v2

@@ Abstract
With technological advancements [4, 6, 7, 21, 22, 23], AI algorithms has shown great potential in improving the efficiency of echocardiography. For example, Ouyang et al. [14] proposed a video-based deep learning algorithm, i.e., EchoNet-Dynamic, that can automatically make accurate assessments of cardiac function. More importantly, with the rise of large language models [2, 16, 17, 20] and multi-modal learning [8, 10, 11, 15], the interpretation of echocardiograms [3] has shown increasing improvement, demonstrating excellent performance in tasks such as pulmonary artery pressure estimation, left ventricular hypertrophy, heart failure, and left atrial enlargement. These AI-assisted diagnostic tools have demonstrated performance almost comparable to human experts. However, this comes with a prerequisite: the acquisition of high-quality echocardiograms. In regions with scarce medical resources, there are often no sonographers available to obtain high-quality echocardiograms. In such cases, the powerful capabilities of these AI-assisted diagnostic tools cannot be fully utilized. Few works [5, 9, 13] have focused on how to use AI technology to assist inexperienced sonographers in accurately acquiring target planes. Recently, Jiang et al. [9] proposed a cardiac dreamer, which only focuses on supervised learning of the 3D-structure of the heart and serves as a "heart map" for the probe guidance task. This work demonstrates a potential AI-assisted scanning method that is expected to improve the scanning skills of novices.
@@
According to the clinical experience of sonographers, understanding both the 2D and 3D structures is crucial for efficient scanning. For example, when you need to adjust the probe’s viewing angle on a particular plane to capture specific anatomical structures, you must have a good understanding of the spatial positions of those anatomical structures on that plane (Fig.1 Left). Undoubtedly, when navigating between different planes, it is essential for sonographers to grasp the spatial relationships within the three-dimensional space (Fig.1 Right). Thus, in this paper, we propose a 2D-3D joint structure-aware pre-training framework to obtain a data-driven cardiac world model that benefits ultrasound scanning. Specifically, we require the world model to learn important spatial relationships in the following ways: (1) for understanding two-dimensional structures, we use a masking approach that requires the world model to predict features at adjacent positions in the two-dimensional plane; (2) for understanding three-dimensional structures, we provide information on the positional changes of two planes in 3D space, requiring the world model to predict the features of the target plane after the positional change. We further collected expert operational data on acquiring the most common ten standard planes from 364 routine clinical scans, resulting in 1.36 million sample pairs gathered by three certified sonographers, to enhance the model’s learning of generalizable cardiac structure knowledge. Results on the downstream probe guidance task indicate that the proposed pre-training method learns useful knowledge for assisting in the acquisition of echocardiograms.
@@ 2.1 Preliminary
Context and Prediction. For the context features, patches not belonging to the target blocks are randomly selected and input into the context encoder to obtain features . Subsequently, the predictor (world model) generates the features of target blocks based on the context features and positional embeddings which indicate the location of context and target blocks. We denote the predicted target blocks’ feature as .
@@
Echocardiography involves acquiring high-quality echocardiogram and subsequently conducting analysis and diagnosis based on these images. However, previous researches [3, 14] have primarily focused on understanding and analyzing the two-dimensional plane, neglecting how AI can assit in acquiring high-quality echocardiogram. While understanding the 2D structure of a plane can aid in scanning, especially when some features of the standard view are already visible, this alone is insufficient. Particularly when transitioning from one standard view to another, a deep understanding of the heart’s 3D structure is essential. Therefore, we propose a 2D-3D joint structure-aware pre-training framework, as illustrated in Fig.2. The core insight is predicting the visual features of structures at target locations based on given 2D and 3D positional conditions, thereby learning the mapping relationship between spatial positions and visual features. Next, we provide details of the modeling and training method.
@@ 3.1 Implementation Details
Downstream task. For the probe guidance task, we adopted the framework and procedure proposed in [9]. The input for this task is an ultrasound image, and the output is the probe position adjustment needed to achieve a specific standard view. This task aims to assist junior ultrasound medical personnel in scanning, enhancing the success rate and quality of view acquisition. During fine-tuning, our pre-trained world model was loaded and optimized for 5 epochs with a batch size of 1024 on 8 Nvidia RTX-4090 GPUs. The learning rate was set to 1e-4, using a cosine scheduler with a final learning rate of 1e-6. The optimizer was set to AdamW.
@@ 4 Conclusion and Discussion
In this work, we propose a 2D-3D joint structure-aware pre-training framework to enhance the cardiac world model’s understanding of spatial relationships within two-dimensional structures on a single view and the three-dimensional spatial relationships between different views. We innovatively designed a self-supervised learning task that predicts visual features based on both two-dimensional and three-dimensional spatial information. To support large-scale self-supervised pre-training, we collected over a million ultrasound image and 3D pose data pairs. After pre-training on the large-scale dataset, considerable improvement were observed in downstream probe guidance tasks across the ten standard views. In the future, we will attempt to: (1) validate our pre-trained model in more downstream tasks that require a comprehensive understanding of the heart’s 2D and 3D structures; (2) validate our pre-trained model in real clinical settings, aiming to directly translate algorithmic improvements into enhanced medical outcomes or increased efficiency.
@@
This work was supported in part by the National Key R&D Program of China (2021ZD0140407), the NSFC (62321005) and the Deng Feng Fund.
4.0.2 Disclosure of Interests.
The authors have no competing interests to declare that are relevant to the content of this article.